Compute the standard deviation σ for ages of British nurses in 1851. Assume that the table below shows the age distribution of nurses in Great Britain in 1851. Round your answer to nearest hundredth.

Age range (yr)

20–29

30–39

40–49

50–59

60–69

70–79

80+

Midpoint (x)

24.5

34.5

44.5

54.5

64.5

75.5

84.5

Percent of nurses

5.8%

9.7%

19.4%

29.1%

24.9%

9.0%

2.1%
Answer #1

The table for calculations will be:

24.5 34.5 44.5 54.5 64.5 74.5 84.5 р(х) 0.058 0.097 0.194 0.291 0.249 0.09 0.021 x*p(x) p(x)(x Mean)A2 1.421 3.3465 8.633 15.8595 16.0605 6.705 1.7745 49.7924 36.1315 16.7791 0.1426 28.5080 38.5641 19.7923 Sum 53.8 189.7100

Hence,

\sigma = \sqrt{\sum p(x)(x- Mean)^2} = \sqrt{189.71} = 13.77